Show that 1.0 mol of ideal gas under STP (standard temperature of exactly 0 oC and pressure of exactly 1.0



atm or 101.3 kPa) occupies a volume of 22.4 L.

1
Expert's answer
2022-03-15T10:35:34-0400

Solution.

"\nu=1 mol;"

"T=273K;"

"P=101300Pa;"

"PV=\nu RT\implies V=\dfrac{\nu RT}{P};"

"V=\dfrac{1\sdot8.31\sdot273}{101300}=0.0224m^3=22.4L;"

Answer: "V=22.4L;"